OpenStudy (anonymous):

Can somebody teach me how to solve a problem like this: x+3/4 = 7/8

OpenStudy (anonymous):

x + 3/4 = 7/8 x = 7/8 - 3/4 = 7/8 - 6/8 = 1/8

OpenStudy (amistre64):

multiply it all by 4; then by 2 then divide

OpenStudy (amistre64):

subtract and divide .... forgot a step

OpenStudy (amistre64):

x+3/4 = 7/8 4x+3 = 7/2 8x+6 = 7 8x = 1 x = 1/8